Lauren's age is 11 years more than four times Robert's age. If the difference of the smaller age from the larger age is 2, how old is Lauren?
Answer by partha_ban(41)

If Lauren's age is 11 years more than four times Robert's age, their age difference cannot be 2. Assuming age difference 20, the solution will be as below:
Let Robert's age be r
Therefore Lauren's age = 4*r + 11
By condition, 4r + 11 - r = 20
or, 4r - r = 20 - 11
or, 3r = 9
or, r = 9/3 = 3
Therefore Lauren's age = 4*3+11 = 12 + 11 = 23 years